Consider a labor market in which LD = 400 – 4W and LS = 250 + 3W. Provide a properly labeled and appropriately articulated Labor Market Model given the labor supply and demand equations provided.
Answer #1

Labor demand (LD):

Given,

LD = Q = 400 – 4W

Now if (Q = 0); 0 = 400 – 4W; 4W = 400; W = 100. The coordinate (Q, W) is (0, 100).

Now if (W = 0); Q = 400 – 0; Q = 400. The coordinate (Q, W) is (400, 0).

Labor supply (LS):

Given,

LS = Q = 250 + 3W

Now if (Q = 0); 0 = 250 + 3W; 3W = - 250; W = -83. The coordinate (Q, W) is (0, -83).

Equilibrium:

LD = LS

400 – 4W = 250 + 3W

7W = 150

W = 150/7 = 21

Now, by putting this value in LD or LS

LD = Q = 400 – 4W

Q = 400 – 4 × 21

    = 400 – 84

    = 316

The coordinate (Q, W) is (316, 21).